# Hybrid Deep Learning Model to Estimate Cognitive Effort from fNIRS Signals in Educational Game Playing

### v2

@@ Abstract.
This study estimates cognitive effort (CE) based on functional near- infrared spectroscopy (fNIRS) data and performance scores using a hybrid deep learning (DeepNet) model. The estimation of CE enables educators to modify material to enhance learning effectiveness and student engagement. In this study, we collected oxygenated hemoglobin $\Delta HbO$ using fNIRS during an educational quiz game. Participants (n=16) responded to 16 questions in a Unity-based educational game, each within a 30-second response time limit. The quiz was divided into two sessions, each session consisting of two task segments. Each segment had four questions. There was a short rest period between the two segments and a longer rest period after each session. We used machine learning and deep learning models to predict the performance score from the fNIRS signal, $\Delta HbO$ and compared traditional machine learning and DeepNet models to determine which approach provides better accuracy in predicting performance scores. The result shows that the GRU gives better performance with 67% than other models. After the prediction, we used the predicted score and the $\Delta HbO$ to observe cognitive effort by calculating relative neural efficiency and involvement in our test cases. Our result shows that even with moderate accuracy, the predicted CE closely follow the actual trends. This findings can be helpful in designing and improving learning environments and provide valuable insights into learning materials.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: XXXXX; XX; isbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)ccs: Applied computing Interactive learning environmentsccs: Applied computing Learning management systemsccs: Applied computing Computer-assisted instruction
Figure 1. The overview of the proposed system. (a) data collection: educational quiz gameplay was used to measure cognitive effort (CE) while recording fNIRS signals. (b) pre-processing: signals underwent pre-processing including filtering and $\Delta HbO$ extraction, and (c) models: various models including a hybrid deep learning model (CNN-GRU) was trained to predict performance scores from the signals(d) output: using the predicted scores and $\Delta HbO$, we estimated CE via Relative Neural Efficiency (RNE) and Involvement (RNI).
[…]
This teaser figure has four blocks showing the overall pipeline. In the leftmost picture (a), the data collection procedure has been showed. The image shows that the participant played an educational game, and we collected the score and oxygenated hemoglobin ($\Delta HbO$). We used an fNIRS device to collect $\Delta HbO$ and saved the score. Then, we combined all the data and created a dataset. The next part shows the steps of pre-processing (b). $\Delta HbO$ signals underwent pre-processing including filtering and $\Delta HbO$ extraction. (c) We showed that the dataset was sent to the models, and after that, we predicted the score (d). Then, based on the $\Delta HbO$ and score, we estimated the cognitive effort.
@@ 1. Introduction
Cognitive effort (CE) refers to the level of cognitive load an individual experiences while learning. This is associated with increased neural activity during problem-solving, decision-making, and learning. In recent years, understanding CE in education-based studies has increased (Getchell and Shewokis, 2023; Benerradi et al., 2019; Sharmin et al., 2024a). When students are learning new or complex tasks, measuring CE helps us understand how hard their brains work and how to improve the design of the educational content. Relative Neural Efficiency (RNE) and Relative Neural Involvement (RNI) are two metrics that can be used to explain CE with behavioral performance  (Getchell and Shewokis, 2023; Koiler et al., 2022; Reddy et al., 2022; Sharmin et al., 2024a). These metrics can be measured using hemodynamic response collected from the prefrontal cortex (PFC) using functional Near-Infrared Spectroscopy (fNIRS) and performance score during a task. The PFC is responsible for executive functions such as decision-making, memory retrieval, and cognitive flexibility, which are critical in educational contexts. RNE indicates how efficiently an individual processes information and performs a task with minimal CE. Higher RNE indicates low effort and high performance, and lower RNE suggests high effort and low performance. On the other hand, RNI measures motivation and engagement. High RNI indicates the user is more engaged and motivated during a task (Paas et al., 2003; Shewokis et al., 2015; Getchell and Shewokis, 2023; Koiler et al., 2022; Sharmin et al., 2024a) Thus, understanding RNE and RNI may help to optimize learning strategies, improve engagement, and enhance cognitive performance.
Traditional methods of evaluating student efficiency and involvement rely on test assessments, self-reports, or reaction times, which are often subjective and inconsistent (De La Cruz et al., 2022; Sharmin et al., 2024a, b). Also, a few studies measured students’ cognitive load, engagement, task difficulty, and efficiency based on only neuron activities(Pan et al., 2020; Sharmin and Barmaki, 2024; Sharmin et al., 2024b). Previously, different AI techniques have been applied in the field of brain signal analysis, such as cognitive workload estimation (Cao et al., 2022), pain level detection (Fernandez Rojas et al., 2019), brain-computer interfaces (BCI) controlling (Hamid et al., 2022), and mental health diagnosis (Kim et al., 2023b). These algorithms are capable of extracting complex patterns from neurophysiological data (Lu et al., 2020). A few recent studies used hemodynamic responses to understand learning efficiency along with subjective measures (Pan et al., 2020; Jeun et al., 2022; Oku and Sato, 2021). However, CE prediction using fNIRS and machine or DeepNet remains an open research challenge.
@@
By applying a Convolutional Neural Network- Gated Recurrent Unit (CNN-GRU) DeepNet model to estimate performance from fNIRS signals, we evaluate the possibility of predicting RNE and RNI, allowing the detection of cognitive fatigue, efficient learning, and recovery effects. To visualize, we divided the cartesian plot into four sub-parts. The High Effort + High Involvement (HE+HI) cognitive state is a critical learning phase observed during skill acquisition, especially in complex or multitasking environments. Individuals in the HE+HI quadrant exert substantial cognitive effort while remaining highly engaged with the task. This state often occurs in the early stages of training, where performance is still developing and mental workload is high due to unfamiliarity with the task or high task demands (Reddy et al., 2022). This quadrant should not be viewed negatively but rather as a productive struggle stage. In this phase, learners are cognitively investing in understanding the task structure, strategies, or switching between sub-tasks (Reddy et al., 2022). Over time, with sufficient practice, effective learners typically transition from HE+HI to LE+HI (Low Effort + High Involvement), showing improved neural efficiency while maintaining motivation (Reddy et al., 2022). Participants in the HE+HI quadrant often exhibit activation in task-relevant prefrontal areas, which supports task switching, attention, and working memory—further validating the cognitive demands being handled during this phase.
This study used hemodynamic responses collected by an fNIRS device to predict performance scores and measure CE by determining RNE and RNI in an educational game setting. Sixteen participants answered 16 questions in two quiz sessions, with fNIRS data recorded at a sampling rate of 10 Hz. We compare the effectiveness of various DeepNet models, such as CNN, LSTM, BiLSTM, and a hybrid CNN-GRU model for quiz score prediction. Then, we used these predicted scores and $\Delta HbO$ to estimate RNE and RNI.
@@
  • RQ1: How does proposed CNN-GRU model compare to traditional ML approaches in predicting performance using brain signals?
  • RQ2 Can we determine the brain region most responsible for predicting cognitive effort using deep learning interpretability techniques on fNIRS data?
  • RQ3: Can a DeepNet model trained on fNIRS-derived features predict performance scores that meaningfully reflect cognitive effort through derived RNE and RNI metrics?
@@
While previous studies have applied machine learning and DeepNet models for cognitive workload estimation, they have primarily focused on classification tasks using standard methods such as n-back. These approaches often rely on subject-specific training and do not predict performance continuously, limiting their applicability in real-time learning environments. In contrast, our work introduces a novel dataset collected during an educational quiz game, capturing both fNIRS hemodynamic responses and performance scores. We predict continuous performance and estimate cognitive effort using relative neural efficiency (RNE) and involvement (RNI). Our hybrid CNN-GRU DeepNet model outperforms both traditional machine learning and standalone DL architectures. Furthermore, we demonstrate that GRU-derived features generalize well when used with XGBoost which suggests strong potential for cross-user adaptability.
@@
The paper is organized as follows. section 2 reviews relevant literature on cognitive effort estimation, highlighting existing methods and identifying research gaps. section 3 outlines our experimental methodology, detailing participants, apparatus, study procedure, cognitive effort measurement methods, data acquisition procedures, and pre-processing techniques. In section 4, we present the proposed hybrid CNN-GRU DeepNet model architecture to evaluate the models. section 5 provides results comparing various machine learning and DeepNet models and analyzes predicted versus actual CE using RNE and RNI. section 6 discusses findings about the research questions, addressing practical implications, limitations, and possible improvements. Finally, section 7 summarizes the study’s findings and highlights its educational contributions.
2. Related Work
Hemodynamic responses using fNIRS have been used widely in cognitive workload assessment and improvement of controlling BCI applications. This is also used in neuro-degenerative disease studies such as Alzheimer’s, Attention-deficit/hyperactivity disorder (ADHD), autism, depression, etc. Machine learning (ML) and DeepNet (DL) have been applied in analyzing fNIRS data across various domains. Traditional ML models such as support vector machine (SVM) (Khan et al., 2021), k-nearest neighbors algorithm (KNN) (De La Cruz et al., 2022), linear discriminant analysis (LDA) (Grässler et al., 2021), etc., have been used for classification tasks, while DeepNet architectures like CNN and LSTM (Lu et al., 2020) have demonstrated superior accuracy in decoding brain activity. Researchers have also leveraged generative models (CNN-GANs) to address data scarcity issues and improve single-trial classification accuracy (Zhang et al., 2023). These advancements highlight the increasing role of ML/DL in fNIRS-based cognitive assessment that forms the foundation for applications in workload estimation and performance prediction.
fNIRS and Machine Learning in Brain Signal Analysis:
@@ 2. Related Work
Deep and machine learning models are being applied to hemodynamic responses collected by fNIRS to control BCI applications by decoding speech (Cooney et al., 2022), motor imagery (Chiarelli et al., 2018; Ma et al., 2021), and finger tapping or movement (Khan et al., 2021; Ortega and Faisal, 2021; Jabbar et al., 2020). These algorithms are also used in health sector to classify pain (Fernandez Rojas et al., 2024; Eken et al., 2024), various neuro-degenerative diseases such as Alzheimer’s (Cicalese et al., 2020; Kim et al., 2023a), Parkinson’s (Quattrocelli et al., 2024; Guevara et al., 2024), ADHD (Shin et al., 2023; Zhao et al., 2021), depression (Li et al., 2025) etc. Detection of fatigue (Tanveer et al., 2019), cognitive load (Zhang et al., 2024), and cognitive tasks such as arithmetic task classification (Wickramaratne and Mahmud, 2021; Zhang et al., 2023) from fNIRS data using ML models gives satisfying results. DeepNet models, particularly CNN and long short-term memory (LSTM), have been effective in decoding brain activity from fNIRS signals with high classification accuracies. Wickramaratne and Mahmud proposed a CNN-based model for mental arithmetic classification and got 87.14% accuracy (Wickramaratne and Mahmud, 2021). Zhang et al. showed in their work that the CGAN-rIRN approach improves the single-trial accuracy (92.19%) for mental arithmetic and mental singing tasks in both the data augmentation and classifier (Zhang et al., 2023). Other models, such as the LSTM model for binary decision-making in BCI, achieved 83.89% accuracy (Erdoğan et al., 2024). Some studies have explored the integration of electroencephalogram (EEG) and fNIRS for improved performance in BCI applications (Cooney et al., 2022; Ortega and Faisal, 2021) by decoding speech (accuracy 87.18%) and moving hands (accuracy 54.4%) using DeepNet methods. EEG and fNIRS were also used in the health sector to classify Alzheimer’s disease (Cicalese et al., 2020), and mild cognitive impairment (Grässler et al., 2021) using LDA. These data were also used for working memory load classification using SVM (accuracy 100%) (Khan et al., 2021) and KNN (accuracy 72%) (De La Cruz et al., 2022).
Beyond these applications, fNIRS has been increasingly applied to predict cognitive workload, fatigue, and performance efficiency, as discussed in the next section.
fNIRS for Cognitive Workload, Fatigue, and Performance Prediction:
Beyond application in BCI and health sectors, researchers have increasingly explored how fNIRS-based models can predict cognitive workload and performance in real-world settings. fNIRS has been increasingly applied to predict cognitive workload and performance. Zhang et al. showed that DeepNet models can predict pilot workload in flight simulations using CNN-attention mechanisms with 88.69% accuracy (Zhang et al., 2024). Similarly, Grimaldi et al. proposed a CNN-LSTM model to predict perceptual load (Grimaldi et al., 2024a) and cognitive load (Grimaldi et al., 2024b) forecasting in piloting tasks. They stated that an LSTM-based approach demonstrated superior performance in short-term workload prediction (Grimaldi et al., 2024b). In another study, Gado et al. classified cognitive workload in NASA task-load (NASA-TLX) simulations using logistic regression, achieving 92.4% accuracy (Gado et al., 2023). Fatigue has also been investigated using fNIRS signals (Ma et al., 2025; Gado et al., 2023; Haroon et al., 2024; Varandas et al., 2022). For instance, fNIRS-based models were employed to classify cognitive fatigue during a digital learning task, achieving an accuracy of 70.91% ± 13.67% using Random Forest (RF)(Varandas et al., 2022). A study combining physiological signals with fNIRS for cognitive fatigue detection using RF reported 94.6% accuracy (Haroon et al., 2024). Ma et al. applied cnn-rESnET to detect fatigue and got 97.78% accuracy in their task (Ma et al., 2025). Most studies have focused on binary or categorical workload classification rather than continuous performance prediction.
In addition to workload and fatigue assessment, recent studies have explored the use of ML/DL for performance efficiency estimation, particularly in educational and learning environments. The following section discusses related works in this domain.
fNIRS in Performance and Efficiency Prediction:
While most studies focus on classifying cognitive workload, researchers have also begun investigating how ML/DL can predict performance efficiency and engagement based on fNIRS signals by monitoring neural activity in different learning and task-based environments.
Jeun et al. used logistic regression to predict task difficulty based on PFC activity before training. The model achieved 86% accuracy in predicting the appropriate difficulty level. After three weeks of training, participants showed improved executive function and reduced PFC activation, indicating increased neural efficiency (Jeun et al., 2022). However, the study was limited by a small sample size and lacked DeepNet approaches for improved accuracy.
Pan et al. applied ML techniques to classify instructional approaches based on brain-to-brain coupling between instructors and learners. They found that machine learning models could distinguish between scaffolding-based learning and explanation-based learning better than single-brain models that highlights the role of interactive engagement in learning efficiency (Pan et al., 2020).
Another approach was proposed by Saikia, where k-means clustering has been used to classify individuals using fNIRS data during a working memory task. Participants performed n-back tasks at four difficulty levels, and clustering revealed three distinct response groups. The study also introduced inverse efficiency score (IES) as a refined performance metric that provided clearer distinctions between difficulty levels compared to traditional methods (Saikia, 2023). However, the model faced challenges in generalizability due to individual differences in brain responses.
Oku and Sato used fNIRS data to model with random forests and logistic regression to classify student engagement and predict quiz performance in online learning platforms. The random forest model achieved a ROC-AUC of 0.67, while logistic regression achieved 0.65, identifying specific brain regions (F4-F6 and AF3-AFz) as key predictors of performance (Oku and Sato, 2021).
@@ 3.1. User Study
Figure 2. Overview of the study procedures: Participants first received an introduction to the study and completed a demographic questionnaire. A training session was provided to familiarize them with the experimental interface and task format. The main session began with a 20-second resting baseline, followed by the first task segment consisting of four questions. After a short 20-second rest, the second segment commenced. This was followed by a longer break and the administration of post-task questionnaires such as the post-test and user experience (UX). The same procedure was repeated for a second session after 10-12 minutes break.
[…]
Figure 2: Overview of the study procedure. Overview of the experimental protocol. Participants first received an introduction to the study and completed a demographic questionnaire. A training session was provided to familiarize them with the experimental interface and task format. The main session began with a 20-second resting baseline, followed by the first task segment consisting of four questions. After a short 20-second rest, the second segment commenced. This was followed by a longer break and administration of post-task questionnaires. The same procedure was repeated for a second session after 10-12 minutes break.The entire protocol was designed to assess participant responses across structured task blocks interspersed with rest periods.
Figure 2 represents the steps of the user study. The participants used a unity-based educational game. The game was designed to evaluate participants’ understanding of the basic structure of graph concepts, such as nodes, edges, loops, etc. Throughout the study, we recorded the participants’ quiz scores and hemodynamic responses from the fNIRS. The Qualtrics survey platform was used to collect responses to demographic questionnaire and pre and post-test questions. The study was approved by the institutional review board.
3.1.1. Apparatus
Figure 3. (a) Experimental setup: Participant playing an educational game wearing fNIRS headband. This setup captures hemodynamic responses in the prefrontal cortex, with data acquisition managed through COBI software on a secondary laptop. (b) 16 channels used for collecting $\Delta HbO$
[…]
Figure 3: (a) Experimental setup: Participant playing an educational game on one laptop wearing fNIRS headband. The fNIRS device is attached to another computer, which captures hemodynamic responses in the prefrontal cortex, with data acquisition managed through COBI software on a secondary laptop. (b) We can see an image of a brain. and pointing 16 channel of pre-frontal cortex that we used for collecting $\Delta HbO$
Figure 3 presents study setup and fNIRS headband orientation. We used two computers and an fNIRS device (Figure 3 (a)). We used an 18-channel fNIRS imager 2000S (fNIRS Device LLC, Potomac, MD, USA), 4 light emitters, wavelength ($\lambda$) 730-850 nm with average power $<1$ mW) sampled at 10 Hz. The detectors were separated by 2.5 cm with a penetrating depth of approximately 1.2 cm. This headband was placed on participants’ foreheads, and fNIRS collected data for 16 channels (Figure 3(b)). The user played the educational quiz games on a Dell XPS touchscreen laptop (L1) [Intel (R) Core(TM) i7-7700HQ processor at 2.80 GHz]. Another desktop computer (L2) [Intel(R) Core(TM) i7-10700T processor at 2.00 GHz] was connected to the fNIRS device. The laptop was used to collect fNIRS data using Cognitive Optical Brain Imaging (COBI) Studio Software. We also used fNIRSoft Software (Version 4.9) to apply filters on $\Delta HbO$ data and generate .CSV file for further processing.
3.1.2. Study Procedure
After signing the consent form and filling out the demographic questionnaires, the participants first experienced a demo version where the study was clearly explained and instructions about sitting still as much as possible (see  Figure 2).
After appearing on a ten-question pre-test, they watched the recorded video tutorial where they learned the basic terminologies of the graph. In our study, we divided our 16 questions into two sessions, with each session having two segments and each segment having four questions. Participants had up to 30 seconds to answer each question, and after each question, they had five seconds of feedback. So, each segment was 140 seconds $((4\times 30=120+4\times 5=20)=140)$ long. In between the two segments, there was a 20-second rest period. After the first session, they had a 6-10 minute break. After each session, the participants completed a post-test on the same graph topic. We counterbalanced the session sequence of the questions to mitigate learning curves.

### v3

@@ Abstract.
This study estimates cognitive effort based on functional near-infrared spectroscopy data and performance scores using a hybrid DeepNet model. The estimation of cognitive effort enables educators to modify material to enhance learning effectiveness and student engagement. In this study, we collected oxygenated hemoglobin using functional near-infrared spectroscopy during an educational quiz game. Participants (n=16) responded to 16 questions in a Unity-based educational game, each within a 30-second response time limit. We used DeepNet models to predict the performance score from the oxygenated hemoglobin, and compared traditional machine learning and DeepNet models to determine which approach provides better accuracy in predicting performance scores. The result shows that the proposed CNN-GRU gives better performance with 73% than other models. After the prediction, we used the predicted score and the oxygenated hemoglobin to observe cognitive effort by calculating relative neural efficiency and involvement in our test cases. Our result shows that even with moderate accuracy, the predicted cognitive effort closely follow the actual trends. This findings can be helpful in designing and improving learning environments and provide valuable insights into learning materials.
@@
journalyear: 2025copyright: ccconference: Companion Proceedings of the 27th International Conference on Multimodal Interaction; October 13–17, 2025; Canberra, ACT, Australiabooktitle: Companion Proceedings of the 27th International Conference onMultimodal Interaction (ICMI Companion ’25), October 13–17, 2025, Canberra, ACT, Australiadoi: 10.1145/3747327.3764901isbn: 979-8-4007-2076-5/2025/10ccs: Human-centered computing Human computer interaction (HCI)$c$s: Applied computing Interactive learning environmentsccs: Applied computing Learning management systemsccs: Applied computing Computer-assisted instruction
Figure 1. The overview of the proposed system. (a) Data collection: educational quiz as learning material was used to measure cognitive effort (CE) while recording fNIRS signals. (b) Pre-processing: signals underwent pre-processing including filtering and oxygenated hemoglobin ($\Delta HbO$) extraction, Balancing was applied only to the training set, not the test set, to preserve evaluation on the original distribution. (c) Models: various models including a hybrid deep learning model (CNN-GRU) was trained to predict performance scores from the signals(d) Output: using the predicted scores and $\Delta HbO$, we estimated CE:
@@ 1. Introduction
Cognitive effort (CE) shows how much mental effort a person uses while learning (Getchell and Shewokis, 2023; Koiler et al., 2022). Cognitive load tells us how active the brain is during a task (Getchell and Shewokis, 2023; Shewokis et al., 2015; Hasan et al., 2023), but it cannot tell if the person is focused or struggling. High brain activity can happen in both cases (Getchell and Shewokis, 2023; Shewokis et al., 2015; Hasan et al., 2023).
CE solves this problem by looking at both brain activity and performance together (Getchell and Shewokis, 2023; Koiler et al., 2022). Functional Near-Infrared Spectroscopy (fNIRS) is used to collect hemodynamic responses from the prefrontal cortex (PFC) as neural activity for executive functions such as decision-making, memory retrieval, and cognitive flexibility. This neural activity and performance score during a task are used to measure two metrics of cognitive effort Relative Neural Efficiency (RNE) and Relative Neural Involvement (RNI) (Getchell and Shewokis, 2023; Koiler et al., 2022; Reddy et al., 2022). RNE shows how efficiently a person completes a task. High RNE means good performance with less effort. Low RNE means more effort and poor results. RNI shows how engaged or motivated a person is (Getchell and Shewokis, 2023; Koiler et al., 2022). A good balance between RNE and RNI means the learner is both focused and efficient.
Old methods like tests and surveys are not always reliable (De La Cruz et al., 2022). Some studies now use brain signals and self-reports to understand learning better (Pan et al., 2020; Jeun et al., 2022; Oku and Sato, 2021; Sharmin et al., 2024c, b). Using fNIRS and machine learning to predict CE is a new and growing research area.
@@
In our previous work we applied various machine learning models to predict score (Sharmin and Barmaki, 2025). In this work, by applying a Convolutional Neural Network- Gated Recurrent Unit (CNN-GRU) DeepNet model to estimate performance from fNIRS signals, we evaluate the possibility of predicting RNE and RNI, allowing the detection of cognitive fatigue, efficient learning, and recovery effects. This study used hemodynamic responses collected by an fNIRS device to predict performance scores and measure CE by determining RNE and RNI in an educational game setting. Sixteen participants answered 16 questions in two quiz sessions, with fNIRS data recorded at a sampling rate of 10 Hz. We compare the effectiveness of various DeepNet models, such as CNN, LSTM, BiLSTM, and a hybrid CNN-GRU model for quiz score prediction. Then, we used these predicted scores and $\Delta HbO$ to estimate RNE and RNI.
@@
  • $RQ_{1}:$ Can deep learning models effectively predict task performance from fNIRS brain signals?
  • $RQ_{2}:$ Which brain regions contribute most to deep learning model predictions of performance, as revealed by interpretability analysis?
  • $RQ_{3}$: Can a DeepNet model trained on fNIRS-derived features predict performance scores that meaningfully reflect cognitive effort through derived RNE and RNI metrics?
@@
While previous studies have applied machine learning and DeepNet models for cognitive workload estimation, they have primarily focused on classification tasks using standard methods such as n-back. These approaches often rely on subject-specific training and do not predict performance continuously, limiting their applicability in real-time learning environments. In contrast, our work introduces a novel dataset collected during an educational quiz game, capturing both fNIRS hemodynamic responses and performance scores. We predict continuous performance and estimate cognitive effort using RNE and RNI. Our hybrid CNN-GRU DeepNet model outperforms both traditional machine learning and standalone DL architectures. Furthermore, we demonstrate that GRU-derived features generalize well when used with XGBoost which suggests strong potential for cross-user adaptability.
@@
The paper is organized as follows. section 2 reviews relevant literature on fNIRS data analysis, highlighting existing methods and identifying research gaps. section 3 outlines our experimental methodology, detailing participants, apparatus, study procedure, cognitive effort measurement methods, data acquisition procedures, and pre-processing techniques. In section 4, we present the proposed hybrid CNN-GRU DeepNet model architecture to evaluate the models. section 5 provides results comparing various machine learning and DeepNet models and analyzes predicted versus actual CE using RNE and RNI. section 6 discusses findings about the research questions, addressing practical implications, limitations, and possible improvements. Finally, section 7 summarizes the study’s findings and highlights its educational contributions.
2. Related Works
@@
Hemodynamic responses using fNIRS have been widely used in cognitive workload assessment and in improving control of brain-computer interface applications. They have also been applied in studies related to neurodegenerative diseases such as Alzheimer’s, ADHD, autism, and depression (Khan et al., 2021; De La Cruz et al., 2022; Grässler et al., 2021; Cooney et al., 2022; Chiarelli et al., 2018; Ma et al., 2021; Fernandez Rojas et al., 2024; Eken et al., 2024; Cicalese et al., 2020; Kim et al., 2023; Quattrocelli et al., 2024; Guevara et al., 2024; Shin et al., 2023; Zhao et al., 2021; Li et al., 2025). Machine learning and deep learning approaches have been widely applied to fNIRS data across domains, with traditional models such as SVM, KNN, and LDA used for classification tasks (Khan et al., 2021; De La Cruz et al., 2022; Grässler et al., 2021), while deep architectures such as CNN and LSTM have shown higher accuracy in decoding brain activity (Lu et al., 2020). CNN-GAN-based models have been introduced to address data scarcity and improve single-trial classification accuracy (Zhang et al., 2023). These approaches have been successfully applied to decode speech, motor imagery, and finger movement (Cooney et al., 2022; Chiarelli et al., 2018; Ma et al., 2021; Khan et al., 2021; Ortega and Faisal, 2021; Jabbar et al., 2020), and to classify pain, fatigue, and various cognitive and clinical conditions with high accuracy.
CNN and LSTM models have reported robust performance in mental arithmetic and decision-making tasks, with accuracies ranging from 83.89% to 92.19% (Wickramaratne and Mahmud, 2021; Zhang et al., 2023; Erdoğan et al., 2024). Some studies have combined EEG and fNIRS to enhance decoding of speech and hand movement (Cooney et al., 2022; Ortega and Faisal, 2021), and to classify Alzheimer’s and mild cognitive impairment using LDA (Cicalese et al., 2020; Grässler et al., 2021). These models have also achieved strong performance in working memory classification using SVM and KNN (Khan et al., 2021; De La Cruz et al., 2022). Beyond brain-computer interface and clinical use, fNIRS has been increasingly applied to predict cognitive workload, fatigue, and performance. For example, CNN-attention and LSTM models have predicted pilot workload and perceptual load with high accuracy (Zhang et al., 2024; Grimaldi et al., 2024a, b), while logistic regression achieved 92.4% accuracy in NASA-TLX simulations (Gado et al., 2023). Fatigue detection has also shown promising results, with accuracies up to 97.78% using RF and CNN-based models (Ma et al., 2025; Gado et al., 2023; Haroon et al., 2024; Varandas et al., 2022). In the learning domain, logistic regression using pre-training PFC signals predicted task difficulty and neural efficiency with 86% accuracy (Jeun et al., 2022). Brain-to-brain coupling was used to classify instructional strategies, revealing the advantage of scaffolding-based learning (Pan et al., 2020). K-means clustering identified individual brain response groups during n-back tasks and introduced inverse efficiency as a refined performance metric (Saikia, 2023). Quiz performance and engagement have been predicted in online platforms using fNIRS signals, where RF and logistic regression models identified key PFC regions with ROC-AUCs of 0.67 and 0.65 respectively (Oku and Sato, 2021).
While prior studies have successfully predicted mental workload and task-related states from fNIRS signals, fewer have explicitly examined how workload relates to performance outcomes in estimating cognitive effort. In this study, we aim to bridge this gap by jointly analyzing workload indicators and quiz performance to explore their relationship and estimate cognitive effort more comprehensively.
@@ 3.1. User Study
The user study was conducted using a Unity-based educational quiz game designed to assess participants’ understanding of graph theory concepts such as nodes, edges, and loops. Participants’ quiz scores and hemodynamic responses were recorded using a fNIRS system. An fNIRS headband was placed on the participant’s forehead to record brain activity throughout the task. The headband was part of an 18-channel fNIRS 2000S device (fNIRS Device LLC, USA), equipped with four light emitters operating at 730–850 nm wavelengths and average power less than 1 mW. Sixteen optodes (measurement points) were positioned with 2.5 cm separation, capturing signals from the prefrontal cortex at an approximate depth of 1.2 cm.
The participant’s brain signals were recorded in real time using Cognitive Optical Brain Imaging (COBI) Studio software installed on a secondary desktop computer. This computer was connected directly to the fNIRS device, while the laptop was used exclusively for the quiz interface. During the session, raw light intensity data were collected at 10 Hz, and later processed through fNIRSoft (v4.9), which applied filters and converted the data to oxygenated hemoglobin ($\Delta HbO$) values. The entire setup was designed to minimize distraction and allow natural interaction. A detailed description of the study design, participant protocol, and data acquisition setup is available in our previous work (Sharmin et al., 2025; Sharmin and Barmaki, 2025).